1750 kg tow truck comes to take your 950kg car away if the coefficient of friction is 0.420 what force must the truck's engine apply in order to accelerate away at 2.20 m/s^2
Answer #1

Fnet = F_applied - muek*m*g

m*a = F_applied - muek*m*g

==> F_applied = m*a + muek*m*g

= 950*2.2 + 0.42*950*9.8

= 6000.2 N